A carretela pulled by a horse carries a 50.0 kg load that includes a copy of modules distributed to remote barangays. The carretela has a mass of 2.50×10^3 kg. If the carretela moves accelerates at a rate of 5.00 m/s^2, how much pull was exerted by the carabao?


1
Expert's answer
2021-01-04T14:32:06-0500

"F=ma=(2500+50)\cdot5=12750(N)" . Answer